OpenStudy (anonymous):

GIVING MEDAL HELPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPPP for the follow system of equations, what is the x value of the solution? -x+2y=6 6y=x+18

OpenStudy (anonymous):

i need to know how to do this

OpenStudy (anonymous):

we can solve this by substitution. rearrange the top equation so it equals y. -x+2y=6 +x-x+2y=6+x 2y = 6+x y=6/2 + x/2 y=3 + x/2 plug this into the second equation 6y=x+18 6(3 + x/2)=x+18 18 + 6x/2 = x + 18 18 + 3x = x + 18 18 + 3x -18= x + 18 -18 3x = x 3x -x = x -x 2x = 0 x=0